# Uni-AIMS: AI-Powered Microscopy Image Analysis

### v1

@@ 1.3 Opportunities with Machine Learning
In recent years, with the rapid development of machine learning and deep learning techniques in the fields of image recognition and analysis ([redmon2016lookonceunifiedrealtime], [wu2019detectron2], [kirillov2023segany]), microscopy image analysis has encountered new opportunities ([choudhary2022recent]). Data-driven approaches can automatically learn the features of the image, significantly improving the efficiency and accuracy of recognition and analysis. Existing studies have attempted to apply cutting-edge deep learning models to microscopy image analysis, achieving promising results, particularly in medical imaging studies ([MedSAM], [archit2025segment]) and biological cell ([moen2019cellimage]) studies.
@@
Furthermore, compared to the extensive datasets available in medical ([kus2024medsegbench]) and biological fields ([goodman2018datasciencebowl]), publicly accessible high-quality microscopy image datasets in materials science are extremely scarce, failing to meet the data volume requirements for training deep learning models. Additionally, annotation of materials microscopy images demands specialized domain knowledge and is time-consuming and labor-intensive, and the lack of high-quality annotated data further restricts the development and application of intelligent analysis methods.
@@ 1.4 Our Work
Considering the unique characteristics of microscopy images, we propose an instance segmentation model that fully accounts for texture features and morphological diversity. By leveraging advanced deep learning algorithms and architectural innovations, the model is capable of accurately detecting targets across scales and robustly identifying and separating thousands of densely packed instances in cluttered backgrounds.
This research effectively overcomes key challenges in materials microscopy image analysis, including data scarcity, annotation difficulties, and sample diversity, achieving both high efficiency and generalizability across different application domains. The proposed intelligent platform provides a robust automated tool for microstructural analysis across multiple disciplines, aiming to meet the urgent needs of researchers and industrial practitioners for efficient and precise microstructural analysis.
@@ 2 Framework Overview
We have developed a systematic research framework for intelligent recognition and analysis of microscopy images. This framework encompasses data acquisition, construction of a multi-source annotation system, deep learning model training, and multi-task applications, ultimately integrated into a unified intelligent analysis platform. The goal is to address the pressing need for automated micro-structural analysis across various scientific and engineering disciplines. Figure 1 illustrates the core components of this framework and their interconnections.
Figure 1: The proposed microscopy image analysis framework. On the left, microscopy images are acquired from various sources and labeled via a human-in-the-loop approach. In the middle, the labeled data undergoes data enhancement to train the Uni-AIMS foundation model for feature extraction. Below, different analysis tasks and their industry applications are showcased. On the right, our intelligent analysis platform automatically generates reports. Throughout this process, data acquisition, labeling, and model training form a data flywheel that iteratively improves the system.
@@
To facilitate the practical deployment of these research outcomes, we further developed an intelligent analysis platform. This platform supports one-click image processing, batch analysis, high-precision recognition, deep data mining, and visualization. It can automatically generate comprehensive and richly formatted analytical reports, thereby offering a powerful and user-friendly solution for automated and intelligent analysis of microscopy images in both scientific research and industrial applications.
@@ 3.1 Data Engine
In order to support robust instance segmentation model training on high-density electron microscopy (EM) images, we developed a web-based data engine that integrates data management, annotation, and quality control. Unlike natural image segmentation—where there are numerous large-scale publicly available datasets—EM image datasets suffer from severe scarcity due to the rarity of instrumentation, proprietary restrictions, and privacy concerns. Normally, annotation tools (e.g. LabelMe ([wada2018labelme]), Label Studio ([LabelStudio])) fail to scale to our high-density scenarios—2,000 × 2,000 images often contain more than 4,000 individual instances, resulting in unacceptable performance degradation and software instability. To overcome these challenges, we designed and implemented an online platform tailored to our domain, complete with specialized annotation widgets, a multi-stage review pipeline, and iterative model–human collaboration to scale both data and model performance.
@@
To bootstrap the dataset and then scale via model-assisted annotation, we adopted a two-stage iterative approach. In Stage I, domain experts annotate a curated, low-density subset of EM images to produce high-accuracy masks. These gold-standard annotations train an initial instance segmentation model, which then generates pseudo-labels on the remaining unannotated data. In Stage II, a larger pool of trained volunteers refines the pseudo-labels using our annotation tools, and corrected masks are validated through the same quality control pipeline before inclusion. Successive cycles of model retraining and human refinement have expanded our dataset to over 5,200 images.
@@ 3.2 Data Generation
Controlled Generation for paired dataset. LoRA-based fine-tuning solutions enable us to quickly learn the visual characteristics of specific materials on small datasets and produce a large number of similar samples. However, during dataset engine, only the image-mask pairs are meaningful for downstream training. Therefore, to generate paired data, we divide the sample augmentation approach into two phases. In phase one, Image-LoRA is used to perform high-quality fitting of visual concepts of specific materials, while Mask-LoRA is utilized to fit masks of instances of different materials. In subsequent stages, Mask-LoRA and heuristic Python scripts are employed to randomly generate instance segmentation masks in bulk. In phase two, ControlNet ([zhang2023adding]) is introduced, where rendered instance segmentation masks are used as conditional inputs to the diffusion model, with their corresponding original images serving as target outputs. Through training on paired data, the ControlNet enhanced by Image-LoRA can utilize the generated masks from the aforementioned to produce corresponding synthetic images.
@@ 3.3 Dataset
Enhanced by data generation techniques, We assembled a large-scale instance segmentation dataset of 6,342 EM images, containing over 1.2 million annotated instances. The dataset was split into a training set of 5781 images and a test set of 561 images. These images were sourced from two primary channels: (1) commercial customer projects, where industrial partners provided raw EM micrographs under confidentiality agreements; and (2) automated web crawling, employing custom scrapers to collect publicly available EM images subject to license verification and de-duplication. To enrich diversity and leverage existing benchmarks, we incorporated three open public datasets:
  • Cellpose ([stringer2021cellpose]): a generalist cellular segmentation corpus of fluorescence and bright-field micrographs with instance-level labels.
  • EMPS ([yildirim2021bayesian]): a collection of 465 EM images furnished with corresponding instance segmentation maps.
  • First public SEM classification dataset ([aversa2018first]): originally released as the inaugural SEM resource providing image-wise class annotations—but lacking any instance-segmentation labels. we selected 300 particularly challenging images and manually produced instance-level masks for each.
@@
We performed detailed statistical analyses on five key properties of our dataset in Figure 3. In Figure 3a, the distribution of instance counts per image is shown on a logarithmic scale: although the median image contains only 81 instances (red dashed line), the mean is substantially higher ($\approx$237, green dash–dot), and the distribution exhibits a heavy right tail (90th percentile at 498, 95th percentile at 870; blue and purple dotted lines), with a few images containing thousands of instances. Figure 3b displays a log-scale histogram of individual mask areas: most object masks occupy between 10² and 10³ pixels, but both very small and very large instances are present, yielding a positively skewed area distribution. In Figure 3c, a scatter plot of image widths versus heights reveals that image resolutions span a broad range (from roughly 200×300 px up to over 3,000 × 2,000 px), with a dense cluster around 400–1,000 px in each dimension. Then, Figure 3d shows the proportion of images by mask-count bins: approximately 16 % of images contain fewer than 10 masks, 22.7 % contain 11–50, 19.1 % contain 51–100, 17.3 % contain 101–200, and 24.9 % contain more than 200 masks. Finally, Figure 3e illustrates the per-image variation in object instance scale across a representative subset of our dataset. To construct it, we first retained only those images containing more than 200 annotated instances; from this pool we then randomly sampled 100 images. For each selected image we computed the pixel area of every instance and plotted, on a logarithmic scale, the minimum instance area, the interquartile range around the median (blue bar), and the maximum area. The wide vertical spans for most images demonstrate that object scales vary by several orders of magnitude both within individual images and across the dataset as a whole, highlighting the challenge of learning scale-robust detectors. Collectively, these panels highlight pronounced imbalance both in the number and size of instances as well as in image resolutions.
@@ 4.1 Overall Workflow
Our workflow begins with batched microscopy image input, which is followed bytream preprocessing before branching into parallel processes: scale bar recognition and instance segmentation. The instance segmentation module also enables AI-assisted labeling on our own data lalabelingg platform, which helps to refine the quality of training data and further improves algorithm performance. Results from both algorithms feed into metrics calculation, which leads to statistical analysis of the extracted features. The analysis results are visualized and presented through both an interactive GUI (which allows human feedback or correction to recompute metrics) and automated report generation, discussed in Section 5.
@@ 4.2.1 Flow-based Segmentation
Instance segmentation fundamentally addresses the challenge of effectively separating individual instances. Detection-based methods like Mask R-CNN ([he2017mask]) and YOLACT ([bolya2019yolact]) employ a two-stage process of detection followed by segmentation, offering intuitive implementation and clear instance differentiation, but suffering performance degradation with densely packed or heavily overlapping objects. Center-based methods such as StarDist ([schmidt2018]), CenterMask ([Wang_2020_CVPR]) and SOLO ([wang2020solo]) construct masks starting from instance centers, effectively handling convex objects in dense scenarios but performing poorly with irregular shapes. Boundary-aware methods including DCAN ([chen2016dcan]), Deep Watershed ([bai2017deep]) and BASNet ([qin2019basnet]) focus on precise boundary detection between instances, excelling at separating adjacent objects but vulnerable to image noise and blurred boundaries. Flow-based methods like HoverNet ([graham2019hover]), Cellpose ([stringer2021cellpose]), EMPS ([yildirim2021bayesian]), and CellViT ([horst2024cellvit]) predict directional fields from pixels to instance centers, efficiently managing complex shapes in densely packed arrangements—particularly suitable for microscopy image analysis—though requiring computationally intensive post-processing and highly accurate flow prediction.
@@
Our flow segmentation algorithm is illustrated in Figure 5. It integrates local geometric feature extraction with global structure analysis, forming a comprehensive processing framework. The method first extracts local morphological and geometric information from the raw image, including boundary gradient fields and foreground object skeletons. These low-level features accurately capture the morphological structure of objects, laying the foundation for subsequent analysis. When processing closed geometric figures, we apply modified Euler integration to the predicted flow fields, creating pixel dynamics models that guide all foreground pixels toward specific convergence centers. By integrating local geometric information with global flow dynamics, our algorithm achieves precise segmentation of targets with different morphologies and scales, providing a reliable foundation for subsequent quantitative analysis and feature extraction. This method is not only applicable to closed morphological cells and particles but also effectively handles open-structured fiber networks, demonstrating broad application potential.
Figure 5: The overview of our instance segmentation algorithm. The input images are encoded by a common vision encoder to patched image embeddings. The embeddings are optionally enhanced with visual and text prompt tokens and sent into multi-head decoder to get geometry information. After post-processing of pixel dynamic, well-separated geometric objects are extracted.
@@ 4.2.2 Implementation
Our algorithm is independent of specific network architecture, making it possible to switch between neural network models of varying computational complexity based on the specific application scenario. For applications with limited computational resources and no need for interactive prompts, we employ a U-Net structure ([ronneberger2015u]) with 2D convolution and skip connections as the core architecture for encoders and decoders. When sufficient data is available, incorporating larger backbone networks, such as hrnet ([WangSCJDZLMTWLX19]), ViT([dosovitskiy2020image]) or Swin ([liu2021swin]) can further enhance model performance, achieving superior analytical results.
In similar algorithms, post-processing based on Pixel Dynamics often constitutes the main performance bottleneck. To address this issue, we implemented the pixel dynamics processing workflow using CUDA, significantly improving the system’s inference throughput. Currently, our optimization solution can complete comprehensive processing of a single 8K×$\times$8K ultra-high-resolution image in approximately 80 seconds on a single Nvidia 4090D GPU, substantially increasing practical application efficiency.
@@ 4.2.3 Result
The metrics we use to evaluate the segmentation model are average precision (AP). AP is calculated by $\text{AP}=\frac{|TP|}{|TP|+|FP|+|FN|}$ under a certain IoU threshold $T$, where the predicted masks matching true masks over $T$ are regarded as TP. Predicted masks and true masks without any matching are regarded as FP and FN. mAP refers to the average AP over all images.
All compared methods in Table 1 were carefully tuned and fine-tuned on our dataset to ensure a fair comparison. For Cellpose, both the network and the size estimation model were retrained using our data. CP-Unet is a modified U-Net model proposed in the Cellpose method. Despite this, Cellpose exhibits lower mAP scores compared to the other methods. This suboptimal performance can be attributed to the core design of Cellpose, which relies heavily on its size model to estimate the typical object size within an image.
In our dataset, however, there is a large variance in object sizes, and in many cases, multiple scales of objects coexist within the same image. This diversity poses a significant advantage for Uni-AIMS over Cellpose, as Uni-AIMS does not depend on a single object size prior, while Cellpose’s single-size estimation approach fails and leads to reduced segmentation accuracy.
Table 1: Comparison of methods on our SEM dataset.
backbone mAP@0.5 mAP@0.7 mAP@0.9
Cellpose CP-Unet 0.564 0.448 0.247
CellViT ViT-b 0.643 0.538 0.294
Uni-AIMS CP-Unet 0.669 0.586 0.327
hrnet-w40 0.738 0.652 0.378
ViT-b 0.787 0.701 0.417

### v2

@@ 1.3 Opportunities with Machine Learning
In recent years, with the rapid development of machine learning and deep learning techniques in the fields of image recognition and analysis ([redmon2016lookonceunifiedrealtime], [wu2019detectron2], [kirillov2023segany]), microscopy image analysis has encountered new opportunities ([choudhary2022recent]). Data-driven approaches can automatically learn image features, significantly improving the efficiency and accuracy of recognition and analysis. Existing studies have attempted to apply cutting-edge deep learning models to microscopy image analysis, achieving promising results, particularly in medical imaging ([MedSAM], [archit2025segment]) and cellular biology ([moen2019cellimage]).
@@
Furthermore, compared to the extensive datasets available in medical ([kus2024medsegbench]) and biological fields ([goodman2018datasciencebowl]), publicly accessible high-quality microscopy image datasets in materials science are extremely scarce, failing to meet the data volume requirements for training deep learning models. Additionally, annotation of materials microscopy images demands specialized domain knowledge and is both time-consuming and labor-intensive; the lack of high-quality annotated data further restricts the development and application of intelligent analysis methods.
@@ 1.4 Our Work
Considering the unique characteristics of microscopy images, we propose an instance segmentation model that fully accounts for texture features and morphological diversity. By leveraging advanced deep learning algorithms and architectural innovations, the model is capable of accurately detecting targets across scales and robustly identifying and separating thousands of densely packed instances within cluttered backgrounds.
This research effectively overcomes key challenges in materials microscopy image analysis, including data scarcity, annotation difficulties, and sample diversity, achieving both high efficiency and strong generalizability across different application domains. The proposed intelligent platform provides a robust automated tool for microstructural analysis across multiple disciplines, aiming to meet the urgent needs of researchers and industrial practitioners for efficient and precise microstructural analysis.
@@ 2 Framework Overview
We have developed a systematic research framework for intelligent recognition and analysis of microscopy images. This framework encompasses data acquisition, construction of a multi-source annotation system, deep learning model training, and multi-task applications, ultimately integrated into a unified intelligent analysis platform. The goal is to address the pressing need for automated microstructural analysis across various scientific and engineering disciplines. Figure 1 illustrates the core components of this framework and their interconnections.
Figure 1: The proposed microscopy image analysis framework. On the left, microscopy images are acquired from various sources and labeled via a human-in-the-loop approach. In the middle, the labeled data undergoes data augmentation to train the Uni-AIMS foundation model for feature extraction. Below, different analysis tasks and their industry applications are showcased. On the right, our intelligent analysis platform automatically generates reports. Throughout this process, data acquisition, labeling, and model training form a data flywheel that iteratively improves the system.
@@
To facilitate the practical deployment of these research outcomes, we further developed an intelligent analysis platform. This platform supports single-click image processing, batch analysis, high-precision recognition, deep data mining, and visualization. It can automatically generate comprehensive and richly formatted analytical reports, thereby offering a powerful and user-friendly solution for automated and intelligent analysis of microscopy images in both scientific research and industrial applications.
@@ 3.1 Data Engine
In order to support robust instance segmentation model training on high-density electron microscopy (EM) images, we developed a web-based data engine that integrates data management, annotation, and quality control. Unlike natural image segmentation—where there are numerous large-scale publicly available datasets—EM image datasets suffer from severe scarcity due to the rarity of instrumentation, proprietary restrictions, and privacy concerns. Normally, annotation tools (e.g. LabelMe ([wada2018labelme]), Label Studio ([LabelStudio])) fail to scale to our high-density scenarios—2,000 × 2,000 images often contain more than 4,000 individual instances, resulting in unacceptable performance degradation and software instability. To overcome these challenges, we designed and implemented an online platform tailored to our domain, complete with specialized annotation widgets, a multi-stage review pipeline, and iterative model–human collaboration to improve both data scalability and model performance.
@@
To bootstrap the dataset and then scale via model-assisted annotation, we adopted a two-stage iterative approach. In Stage I, domain experts annotate a curated, low-density subset of EM images to produce high-accuracy masks. These gold-standard annotations train an initial instance segmentation model, which then generates pseudo-labels on the remaining unannotated data. In Stage II, a larger pool of trained volunteers refines the pseudo-labels using our annotation tools, and corrected masks are validated through the same quality control pipeline before inclusion. Successive cycles of model retraining and human refinement have expanded our dataset to 5,091 images.
@@ 3.2 Data Generation
Controlled Generation for paired dataset. LoRA-based fine-tuning solutions enable us to quickly learn the visual characteristics of specific materials on small datasets and produce a large number of similar samples. However, within our data engine, only image-mask pairs are meaningful for downstream training. Therefore, to generate paired data, we divide the sample augmentation approach into two phases. In phase one, Image-LoRA is used to perform high-quality fitting of visual concepts of specific materials, while Mask-LoRA is utilized to fit masks of instances of different materials. In subsequent stages, Mask-LoRA and heuristic Python scripts are employed to randomly generate instance segmentation masks in bulk. In phase two, ControlNet ([zhang2023adding]) is introduced, where rendered instance segmentation masks are used as conditional inputs to the diffusion model, with their corresponding original images serving as target outputs. Through training on paired data, the ControlNet enhanced by Image-LoRA can utilize the generated masks from the aforementioned to produce corresponding synthetic images.
@@ 3.3 Dataset
We assembled a large-scale instance segmentation dataset of 5,091 EM images, containing about 3 million annotated instances. We refer to this dataset as UniEM-3M, which has been released and described in detail in our work ([uniem3m]). The dataset was split into a training set of 4128 images and a test set of 963 images. These images were sourced from three primary channels: (1) high-quality, original EMs from academic collaborations, (2) automated web crawling, and (3) the annotation of approximately 500 challenging images from the public dataset ([aversa2018first]).
@@
We performed detailed statistical analyses on five key properties of our dataset in Figure 3. In Figure 3a, the distribution of instance counts per image is shown on a logarithmic scale: although the median image contains 246 instances (red dashed line), the mean is substantially higher ($\approx$586, green dash–dot), and the distribution exhibits a heavy right tail (90th percentile at 1429, 95th percentile at 2342; blue and purple dotted lines), with a few images containing thousands of instances. Figure 3b displays a log-scale histogram of individual mask areas: most object masks occupy between $10^{2}$ and $10^{4}$ pixels, but both very small and very large instances are present, yielding a positively skewed area distribution. In Figure 3c, a scatter plot of image widths versus heights reveals that image resolutions span a broad range (from roughly 200×200 px up to over 8,000 × 6,000 px), with a dense cluster around 400–1,500 px in each dimension. Then, Figure 3d shows the proportion of images by mask-count bins: approximately 20 % of images contain fewer than 10 masks, 11 % contain 51–100, 14 % contain 101–200, 22 % contain 201–500, and 33 % contain more than 500 masks. Finally, Figure 3e illustrates the per-image variation in object instance scale across a representative subset of our dataset. To construct it, we first retained only those images containing more than 200 annotated instances; from this pool we then randomly sampled 100 images. For each selected image we computed the pixel area of every instance and plotted, on a logarithmic scale, the minimum instance area, the interquartile range around the median (blue bar), and the maximum area. The wide vertical spans for most images demonstrate that object scales vary by several orders of magnitude both within individual images and across the dataset as a whole, highlighting the challenge of learning scale-robust detectors. Collectively, these panels highlight pronounced imbalance both in the number and size of instances as well as in image resolutions.
@@ 4.1 Overall Workflow
Our workflow begins with batched microscopy image input, which is followed by image preprocessing before branching into parallel processes: scale bar recognition and instance segmentation. The instance segmentation module also enables AI-assisted labeling on our own data annotation platform, which helps to refine the quality of training data and further improves algorithm performance. Results from both algorithms feed into metrics calculation, which leads to statistical analysis of the extracted features. The analysis results are visualized and presented through both an interactive GUI (which allows human feedback or correction to recompute metrics) and automated report generation, discussed in Section 5.
@@ 4.2.1 Flow-based Segmentation
Instance segmentation fundamentally addresses the challenge of effectively separating individual instances. Detection-based methods like Mask R-CNN ([he2017mask]) employ a two-stage process of detection followed by segmentation, offering intuitive implementation and clear instance differentiation, but suffering performance degradation with densely packed or heavily overlapping objects. In contrast, one-stage methods like YOLACT ([bolya2019yolact]) prioritize speed by producing shared prototype masks, trading off some mask quality. Center-based methods such as CenterMask ([Wang_2020_CVPR]) and SOLO ([wang2020solo]) construct masks starting from instance centers, effectively handling convex objects in dense scenarios but performing poorly with irregular shapes. Universal segmentation frameworks such as Mask2Former ( [cheng2021mask2former]) have been proposed, which leverage a transformer-based decoder with masked attention to unify semantic, instance, and panoptic segmentation. While highly general and flexible, these methods may struggle to maintain fine-grained mask quality in microscopy images, particularly under extreme density or scale variation. StarDist ([schmidt2018]) and its extension CPP-Net ( [chen2023cpp]), although often grouped into center-based approaches, actually model each object as a star-convex polygon via radial distances from a predicted center point, which makes it particularly well-suited for convex cell nuclei but less effective on irregularly shaped structures. Flow-based methods like HoverNet ([graham2019hover]), Cellpose ([stringer2021cellpose]) and CellViT ([horst2024cellvit]) predict directional fields from pixels to instance centers, efficiently managing complex shapes in densely packed arrangements—particularly suitable for microscopy image analysis—though requiring computationally intensive post-processing and highly accurate flow prediction.
@@
Table 1: Baseline performance comparison of various instance segmentation methods on sparse and dense subdatasets.
Class Method Backbone Params Sparse Dense
$mAP@0.5$ $PQ@0.5$ $mAP@0.5$ $PQ@0.5$
Anchor-based Mask R-CNN ResNeXt101 101M 0.542 0.565 - -
Cascade R-CNN ResNeXt101 135M 0.540 0.587 - -
HTC ResNeXt101 137M 0.345 0.461 - -
Anchor-free YOLACT ResNet101 54M 0.307 0.293 - -
Mask2Former Swin-base 107M 0.204 0.064 - -
StarDist SAM-ViT-base 140M 0.236 0.230 0.487 0.438
CPP-Net SAM-ViT-base 140M 0.196 0.187 0.378 0.352
2D Field CellViT SAM-ViT-base 146M 0.589 0.573 0.701 0.604
Cellpose-SAM SAM-ViT-large 304M 0.605 0.633 0.760 0.700
Uni-AIMS SAM-ViT-base 93M 0.824 0.720 0.787 0.703
Our flow-based segmentation algorithm is illustrated in Figure 5. It integrates local geometric feature extraction with global structure analysis, forming a comprehensive processing framework. The method first extracts local morphological and geometric information from the raw image, including boundary gradient fields and foreground object skeletons. These low-level features accurately capture the morphological structure of objects, laying the foundation for subsequent analysis. For closed geometric figures, we apply modified Euler integration to the predicted flow fields, creating pixel dynamics models that guide all foreground pixels toward specific convergence centers. By integrating local geometric information with global flow dynamics, our algorithm achieves precise segmentation of targets with different morphologies and scales, providing a reliable foundation for subsequent quantitative analysis and feature extraction. This method is not only applicable to closed morphological cells and particles but also effectively handles open-structured fiber networks, demonstrating broad application potential.
Figure 5: The overview of our instance segmentation algorithm. The input images are encoded by a common vision encoder to patched image embeddings. The embeddings are optionally enhanced with visual and text prompt tokens and sent into a multi-head decoder to get geometry information. After post-processing of pixel dynamics, well-separated geometric objects are extracted.
@@ 4.2.2 Implementation
Our algorithm is independent of specific network architecture, making it possible to switch between neural network models of varying computational complexity based on the specific application scenario. For applications with limited computational resources and no need for interactive prompts, we employ a U-Net structure ([ronneberger2015u]) with 2D convolution and skip connections as the core architecture for encoders and decoders. When sufficient data is available, incorporating larger backbone networks, such as ResNet ([he2016deep]), ResNeXt ([xie2017aggregated]), ViT ([dosovitskiy2020image]) or Swin ([liu2021swin]) can further enhance model performance, achieving superior analytical results.
In similar algorithms, post-processing based on pixel dynamics often constitutes the main performance bottleneck. To address this issue, we implemented the pixel dynamics processing workflow using CUDA, significantly improving the system’s inference throughput. Currently, our optimization solution can complete comprehensive processing of a single $\times8K$ ultra-high-resolution image in approximately 80 seconds on a single Nvidia 4090D GPU, substantially increasing practical application efficiency.
@@ 4.2.3 Result
We adopt standard metrics for instance segmentation: mean Average Precision (mAP) for mask accuracy, and Panoptic Quality (PQ) ( [kirillov2019panoptic]) at 0.5 for combined segmentation and recognition quality. Following StarDist ( [schmidt2018]), we use AP calculated by $\text{AP}=\frac{|TP|}{|TP|+|FP|+|FN|}$  [^1] In this paper, AP follows the convention in bioimage segmentation [stringer2021cellpose, graham2019hover, schmidt2018], which differs from the COCO-style AP but is widely adopted in microscopy instance segmentation. at an IoU threshold $T$, where the predicted masks matching true masks over $T$ are regarded as TP. Mean AP is the average AP of all images.
Anchor-based methods (e.g., Mask R-CNN, Cascade Mask R-CNN) and recent frameworks such as YOLACT and Mask2Former often suffer from excessive computational cost and degraded accuracy under ultra-dense microscopy scenes. To better reflect these performance discrepancies, we divide our dataset into two subsets: a sparse set (images with instances fewer than 100) and a dense set (more than 100 instances). In contrast, bioimage-oriented methods like Cellpose and HoverNet are generally more robust to both regimes, with memory usage largely independent of the number of instances. All compared methods in Table 1 were carefully tuned and fine-tuned on our dataset to ensure a fair comparison. These experiments demonstrate the superior capability of flow-based methods, particularly Uni-AIMS, in handling high-density microstructures, likely attributable to their effective use of 2D fields for instance separation in crowded scenes.